OpenStudy (spring98):

(7*3*2/7*5)^2*(7^0/5^-3)^3*5^-9=? Can some one help me?

OpenStudy (spring98):

is the answer to this problem 36/25?

OpenStudy (spring98):

@MikuxLuka

OpenStudy (anonymous):

idk dis ;-;

OpenStudy (spring98):

no one is helping me :(

OpenStudy (spring98):

:.(

OpenStudy (spring98):

it's same thing as that. the one i put looks a little confusing but it realy is the same.

OpenStudy (anonymous):

;- ;-; ;-; idk hun

OpenStudy (here_to_help15):

@Spring98 still need help?

OpenStudy (spring98):

yes

OpenStudy (here_to_help15):

\[(\frac{ 7*3*2 }{ 7*5 })^2 *(\frac{ 7^0 }{ 5^{-3} })^3 *5^{-9}=?\]

OpenStudy (here_to_help15):

Thats it correct?

OpenStudy (spring98):

yes

OpenStudy (spring98):

can you help me with this?

OpenStudy (here_to_help15):

Yes just multi tasking is all sorry

OpenStudy (spring98):

what do you mean?

OpenStudy (here_to_help15):

Im doing history essay,English essay,Algebra 2 homework and this ... So doing 4 things at one time is taking a bit.

OpenStudy (spring98):

i know but are you going to help me? because i need this asap

OpenStudy (here_to_help15):

@Jaynator495 @jim_thompson5910 please come so i can finish my own homework :)

OpenStudy (here_to_help15):

@amistre64

jimthompson5910 (jim_thompson5910):

`is the answer to this problem 36/25?` no, the answer is a whole number. Try again @Spring98

jimthompson5910 (jim_thompson5910):

these rules may help http://i.ytimg.com/vi/csQQMqmmcFs/maxresdefault.jpg

OpenStudy (spring98):

is it 252/5/

OpenStudy (spring98):

252/5

OpenStudy (spring98):

is 252/5 the answer?

OpenStudy (here_to_help15):

@Spring98 do you know what whole number is?

OpenStudy (spring98):

no

OpenStudy (spring98):

can you at least help

OpenStudy (spring98):

plz